The length of a rectangle is 5 YARDS more than double the width, and the area of the rectangle is 63 YARDS . Find the dimensions of the rectangle.
Answer by TimothyLamb(4379)

L = 2W + 5
A = LW = 63
L = 63/W
---
L = 2W + 5
63/W = 2W + 5
63 = 2WW + 5W
2WW + 5W - 63 = 0
---
2WW + 5W - 63 = 0
is already in quadratic equation form: ax^2 + bx + c = 0
---
use the quadratic formula to solve for two solutions:
x = ( -b +- sqrt( bb - 4ac ) ) / 2a
---
the two solutions are:
x = 9/2
x = -7
the negative solution doesn't make sense for length, so use the positive solution.
---
W = 9/2
L = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php